Two infinite and parallel sheets of charge have the same surface density. What is the field in the region between the sheets







1
Expert's answer
2021-11-01T19:29:03-0400

We know that

Surface charge density

"\sigma=\frac{Q}{A}"



Now plate left side electric field

"E_L=\frac{-\sigma}{2\epsilon_0}+\frac{\sigma}{2\epsilon_0}=0"

Plate right side electric field

"E_R=\frac{\sigma}{2\epsilon_0}-\frac{\sigma}{2\epsilon_0}=0"

Plate mid point electric field

"E_m=\frac{\sigma}{2\epsilon_0}+\frac{\sigma}{2\epsilon_0}=\frac{\sigma}{\epsilon_0}"

To infinite and parallel sheet between electric field

"E_m=\frac{\sigma}{2\epsilon_0}+\frac{\sigma}{2\epsilon_0}=\frac{\sigma}{\epsilon_0}"